# Safely Learning Controlled Stochastic Dynamics

### v2

@@ A.5 Proof of sample complexity of confidence intervals
We refer the reader to seeger2008information; srinivas2009gaussian; vakili2021information for additional examples and their associated proofs, which show that the growth rate of $\gamma_{N}$ depends on the eigenvalue decay in the Mercer expansion of the kernel, when available.
@@ A.6 Proof of learning rates for system maps estimation at fixed $(\theta,t)$
Following steps 3 and 4 of Theorem 5.3 in bonalli2025non, for $\nu>n/2+\varepsilon$ we establish
@@ A.7 Proof of safe learning of controlled Sobolev dynamics
where $\beta_{N}^{s}\triangleq\lambda^{-1}N^{-1/2}\max_{i\in\llbracket 1,N\rrbracket}|\hat{s}_{\theta_{i},t_{i}}-s(\theta_{i},t_{i})|+\|s\|_{\mathcal{G}}$, $\beta_{N}^{r}\triangleq\lambda^{-1}N^{-1/2}\max_{i\in\llbracket 1,N\rrbracket}|\hat{r}_{\theta_{i},t_{i}}-r(\theta_{i},t_{i})|+\|r\|_{\mathcal{G}}$, and $\beta_{N}^{p}\triangleq\lambda^{-1}N^{-1/2}\max_{i\in\llbracket 1,N\rrbracket}\|\hat{p}_{\theta_{i},t_{i}}(\cdot)-p_{\theta_{i}}(t_{i},\cdot)\|_{L^{\infty}(\mathbb{R}^{n})}+\|p\|_{L^{\infty}(\mathbb{R}^{n};\mathcal{G})}$.

### v3

@@ A.5 Proof of sample complexity of confidence intervals
We refer the reader to Seeger et al. [2008]; Srinivas et al. [2009]; Vakili et al. [2021] for additional examples and their associated proofs, which show that the growth rate of $\gamma_{N}$ depends on the eigenvalue decay in the Mercer expansion of the kernel, when available.
@@ A.6 Proof of learning rates for system maps estimation at fixed $(\theta,t)$
Following steps 3 and 4 of Theorem 5.3 in Bonalli and Rudi [2025], for $\nu>n/2+\varepsilon$ we establish
@@ A.7 Proof of safe learning of controlled Sobolev dynamics
where $\beta_{N}^{s}\triangleq\lambda^{-1/2}\max_{i\in\llbracket 1,N\rrbracket}|\hat{s}_{\theta_{i},t_{i}}-s(\theta_{i},t_{i})|+\|s\|_{\mathcal{G}}$, $\beta_{N}^{r}\triangleq\lambda^{-1/2}\max_{i\in\llbracket 1,N\rrbracket}|\hat{r}_{\theta_{i},t_{i}}-r(\theta_{i},t_{i})|+\|r\|_{\mathcal{G}}$, and $\beta_{N}^{p}\triangleq\lambda^{-1/2}\max_{i\in\llbracket 1,N\rrbracket}\|\hat{p}_{\theta_{i},t_{i}}(\cdot)-p_{\theta_{i}}(t_{i},\cdot)\|_{L^{\infty}(\mathbb{R}^{n})}+\|p\|_{L^{\infty}(\mathbb{R}^{n};\mathcal{G})}$.